Suppose that a ball is drawn at random from a box containing three white & three black balls . After a ball is drawn , it is then replaced & another drawn. What is the probability that of the first four balls drawn, exactly two are white?
1
Expert's answer
2020-06-08T20:51:29-0400

Probability of first drawn white "= \frac{3C2\times3C2}{6C4} = \frac{3}{5}"


Probability of second drawn white with replacement "= \frac{3C2\times3C2}{6C4} = \frac{3}{5}"


Probability of exactly two are white = Probability of first drawn white "\times"

Probability of second drawn white with replacement


Probability of exactly two are white "={\frac{3}{5}}\times\frac {3}{5}= \frac {9}{25}"